### v1

# The Impact of Spectroscopic Redshift Errors on Cosmological Measurements

Abstract
Spectroscopic redshift errors, including redshift uncertainty and catastrophic failures, can bias cosmological measurements from galaxy redshift surveys at sub-percent level. We investigate their impact on full-shape clustering analysis using contaminated mock catalogs. We find that redshift uncertainty introduces a damping effect on the power spectrum. This damping is scale-dependent and absorbed by counterterms in the clustering model, keeping parameter biases below $5\%$ for the DESI survey. Catastrophic failures reduce the power spectrum amplitude by an approximately constant factor scaling with contamination rate $f_{c}$. While this effect is negligible for the DESI ELG populations ($f_{c}=1\%$), the slitless-like errors, combining redshift uncertainty with $f_{c}=5\%$ catastrophics, introduce significant biases in cosmological constraints. For this case, shifts from $6\%$ to $16\%$ ($\sim 2.2\sigma$ level) arise in estimating the fractional growth rate $df\equiv f/f^{\rm{fid}}$ and the log primordial amplitude $\ln(10^{10}A_{s})$. Applying a correction factor $(1-f_{c})^{2}$ on the galaxy power spectrum mitigates the bias but weakens the parameter constraints due to new degeneracies. Alternatively, fixing $f_{c}$ to its expected value during fitting successfully restores the unbiased posterior without loss in constraint. Our results indicate that for space-based slitless surveys such as Euclid, an accurate estimation of $f_{c}$ and its incorporation into the clustering model are essential to get unbiased cosmological constraints. Extending to evolving dark energy and massive neutrino cosmologies, we find that redshift errors do not bias the dark energy properties parametrized by $w_{0}$ and $w_{a}$, but lead to slightly weaker constraints on $\sum m_{\nu}$.
1 Introduction
The large-scale structure (LSS) observed by galaxy redshift surveys has played an important role in advancing our understanding of fundamental physics. Clustering features, such as baryon acoustic oscillations (BAO; [1, 2, 3, 4, 5]), and the full-shape power spectrum [6, 7, 8, 9], have been used to constrain some key quantities in cosmology, including dark energy equation of state $w$, the matter density $\Omega_{\mathrm{m}}$, the structure growth rate $f\sigma_{8}$, neutrino mass $\sum m_{\nu}$ and possible deviation from general relativity. Current leading experiments, the Dark Energy Spectroscopic Instrument [^1] DESI, desi.lbl.gov (DESI; [10, 11]), and the ongoing space telescope Euclid [^2] Euclid, cosmos.esa.int/web/euclid [12, 13] are mapping the 3-dimensional Universe by measuring redshifts for tens to hundreds of millions of galaxies. These surveys will cover over more than 15,000 $\rm deg^{2}$ of the sky, and trace the LSS back to redshift $\sim 3.5$. This large probing volume allows the clustering signal to be measured with minimal cosmic variance, leading to sub-percent precision in cosmological measurements. Given this level of precision, detailed studies on systematic errors are critical to ensure the robustness and reliability of cosmological constraints.
Among various observational systematics, we focus on the impact of spectroscopic redshift errors. These errors arise from inaccuracies in the spectroscopic pipelines [14, 15, 16, 17]. They persist in data with secure redshifts, where calibration and algorithmic processing have already been applied. There are two primary types of errors: redshift uncertainty and redshift catastrophic failures. Redshift uncertainty refers to minor systematic offsets ($\Delta z\sim 10^{-4}$) in all redshift measurements, primarily caused by instrumental noise and the intrinsic width of spectral features. In contrast, redshift catastrophic failures result from the misidentification of spectral lines, leading to severely incorrect redshifts ($\Delta z\sim 10^{-2}$) in a small fraction of data. These systematic errors are essential because they can distort the observed clustering signal and potentially bias the cosmological inference. While previous studies have examined their individual impacts [18, 19, 20, 21, 22, 23], a comprehensive analysis that considers all these effects remains unexplored. This work aims to bridge that gap and explore their combined influence on cosmological inference from full-shape clustering analyses.
The effect of spectroscopic redshift errors can be quantified using repeated redshift observations and incorporated into realistic galaxy mocks (e.g., [24, 25, 26, 21]). By comparing analyses on clean mocks with those contaminated by redshift errors, we can assess their impact on clustering statistics and cosmological constraints. Improved redshift calibration and increased repeat observation samples in DESI [27, 21] provide a better way to model the redshift errors and more precise studies of their impact on cosmological analyses. Additionally, space-based surveys such as Euclid use slitless spectroscopy at near-infrared to measure redshifts [28, 29]. This method requires no pre-selection of targets and is particularly effective for probing high-redshift galaxies. However, it is also more susceptible to redshift errors due to the lower spectrum resolution and the potential overlap of different spectra. These challenges have been addressed in the context of redshift interlopers for Euclid [23]. In this work, we use a hypothetical combined redshift errors to represent the conditions expected in slitless surveys. This approach provides a realistic description of the related systematics and offers practical guidance for interpreting and mitigating their impact in cosmological analyses for slitless survey.
The outline of the paper is as follows: Section˜2 introduces two spectroscopic surveys, DESI and Euclid, and the spectroscopic redshift errors. Section˜3 describes how we construct our contaminated mock catalogs. Section˜4 presents the methodologies used to derive cosmological inferences from the mock data. In Section˜5, we present the impact of redshift errors on clustering statistics and how these propagate into the cosmological parameter constraints, including extensions $w_{0}w_{a}\rm{CDM}$ and $\nu\Lambda\rm{CDM}$ models. Finally, we summarize our findings and conclude in Section˜6.
2 Spectroscopic Survey and Redshift Errors
Spectroscopic surveys measure the redshifts of objects with high precision by analyzing their spectral features, such as emission and absorption lines. Two massive ongoing surveys, DESI and Euclid, are using spectroscopy to measure galaxy clustering with unprecedented accuracy. In its first three years, the ground-based DESI survey has observed over 14 million galaxies [5] and plans to extend its initial plan to 50M galaxies across $18{,}000\,\rm{deg}^{2}$. Its dark matter tracer includes Bright Galaxies (BGS) at $0.1<z<0.4$ [30], Luminous Red Galaxies (LRG) at $0.4<z<1.1$ [31], Emission Line Galaxies (ELG, mainly [O ii] emitters) at $0.8<z<1.6$ [32], and Quasi-Stellar Objects (QSO) at $0.8<z<3.5$ [33]. The Euclid mission conducts a slitless spectroscopic survey from space, aiming to observe tens of millions of galaxies over $15{,}000\,\rm{deg}^{2}$, primarily targeting $\rm{H\,\alpha}$] emitters in the redshift range $0.9<z<1.8$ [13]. Its Quick Data Release 1 is already publicly available [34]. The large cosmic volumes probed by DESI and Euclid enable sub-percent precision in cosmological measurements and require a rigorous understanding of systematic errors, particularly those related to redshift measurements.
In this work, the spectroscopic systematic redshift errors are referred to as redshift errors. The redshift errors exist in both ground-based and space-based galaxy redshift surveys. They include redshift uncertainty, the small statistical fluctuations in redshift measurements, and redshift catastrophic failures, the large errors from the true values. Both types of error persists in secure redshift measurements, i.e., those that have been calibrated and successfully passed the redshift determination pipeline. Consequently, these hidden errors can distort the observed clustering pattern and potentially bias our cosmological analyses. Since both redshift uncertainty and catastrophic failures can be quantified from repeated redshift observations, it is practical to analyze them together. This combined treatment is also relevant for space missions such as Euclid, which apply slitless spectroscopy to measure redshifts and are therefore more susceptible to both types of contamination. In the following, we will describe how redshift uncertainty and redshift catastrophic failures are characterized and modeled, and what are the redshift errors expected in slitless spectroscopic surveys.
2.1 Redshift uncertainty
Redshift uncertainty is an inevitable systematic error in spectroscopic redshift measurements. It arises because the spectral lines used to determine redshifts have specific widths rather than being ideal delta functions. These widths can result from a combination of physical effects (e.g., thermal motions, turbulence, gravitational broadening) and the instrumental resolution. Spectroscopic surveys report redshift uncertainty for different types of targets. For example, in eBOSS galaxy samples, the mean redshift uncertainty for LRG is $65.6\,\rm{km}\cdot s^{-1}$ [35]. For QSO, the typical uncertainty is around 300 $\rm{km}\cdot s^{-1}$ [36]. Repeat observations of ELG indicate that 50% of redshift estimates have a precision better than $20\,\rm{km}\cdot s^{-1}$ [37]. In the recent DESI Data Release 1 (DR1), the redshift precision reaches $\sim 10\,\rm{km}\cdot s^{-1}$ for BGS and ELG, $\sim 50\,\rm{km}\cdot s^{-1}$ for LRG, and $20$-$125\,\rm{km}\cdot s^{-1}$ for QSO in $0.8<z<1.8$ [27].
The effect of redshift uncertainty is equivalent to randomly smearing the peculiar velocity of an observed object along the line-of-sight. This smearing velocity is empirically quantified using redshift difference, $\Delta z$, from repeated observations, given by $\Delta v_{\rm{smear}}=c\,\Delta z/(1+\bar{z})$. The $c$ is the speed of light and $\bar{z}$ is the preferred redshift measurement of the object, typically taken as the mean or the highest-confidence one from repeats. The distribution of observed $\Delta v_{\rm{smear}}$ can be statistically modeled. For galaxies, this distribution is typically Gaussian $\mathcal{N}(\mu,\,\sigma_{\Delta v})$, with $\sigma_{\Delta v}$ characterizing the redshift uncertainty dispersion [35, 25, 26]:
$\mathcal{N}(\Delta v_{\rm{smear}}\,|\,\mu,\,\sigma_{\Delta v})=\frac{1}{\sqrt{2\pi\sigma_{\Delta v}}}\exp\left(-\frac{(\Delta v_{\rm{smear}}-\mu)^{2}}{2\sigma_{\Delta v}}\right)$ (2.1)
For QSO, the distribution of $\Delta v_{\rm{smear}}$ usually exhibits extended wide tails. In such case, a multiple-Gaussian model [36, 33] or a Lorentzian distribution [26], provides a better fit to repeat measurements. In our work, we use the Lorentzian profile, $\mathcal{L}(\,p,\,w_{\Delta v})$, to describe the QSO redshift uncertainty, which can be expressed as:
$\mathcal{L}(\Delta v_{\rm{smear}}\,|\,p,\,w_{\Delta v})=\frac{A}{1+((\Delta v_{\rm{smear}}-p)/w_{\Delta v})^{2}}$ (2.2)
Here, $w_{\Delta v}$ captures the width of the distribution, thus characterizing the redshift uncertainty in the Lorentzian profile. In our analyses, we adopt both Gaussian and Lorentzian profiles to model QSO redshift uncertainty. In this way, we can access how the presence of long-tailed uncertainty distributions may impact clustering and cosmological measurements.
2.2 Redshift catastrophic failures
Redshift catastrophic failures (or redshift catastrophics) arise in a small fraction of redshift measurements, typically identified through significant redshift differences in repeated observations. These errors can result from misidentification of sky residuals or confusion between spectral lines of different galaxies [14, 21, 17]. Faint or ambiguous spectral features can also lead to catastrophically incorrect redshift measurements. The catastrophics contamination rate, $f_{c}$, and the threshold for identifying it vary in different surveys. For example, the WiggleZ dark energy survey reports a 5% $f_{c}$ by defining redshift blunders with $|\Delta z|>400\,\rm{km}\cdot s^{-1}$ (larger than $4\sigma_{\Delta v}$) [14]. In eBOSS, the reported $f_{c}$ is less than 1% for LRGs and ELGs ($|\Delta z|>1000\,\rm{km}\cdot s^{-1}$) [37, 35] and 1.6% for quasars ($|\Delta z|>3000\,\rm{km}\cdot s^{-1}$) [36]. In DESI DR1, the $f_{c}$ is estimated to be less than 1% for all tracers at $z<2.1$ [27].
In the study of DESI DR1 repeated ELG observations, redshift catastrophics are classified into two main categories: sky confusion at $z=1.32$ and random catastrophics [21]. Sky confusion, responsible for 24% of these errors, arises from misidentifying residuals of sky emissions at around 8625 Å and can be approximated as a Gaussian distribution. The remaining 76% random catastrophics encompass various error patterns and are modeled using a log-normal profile as:
$\mathcal{LN}(\Delta\hat{v}_{\rm{catas}}\,|\,\mu_{\text{ran}},\sigma_{\text{ran}}^{2},v_{0})=\frac{A}{\sqrt{2\pi}\sigma_{\text{ran}}\Delta\hat{v}_{\rm{catas}}}\,\exp\left[-\frac{\left(\Delta\hat{v}_{\rm{catas}}-\mu_{\text{ran}}\right)^{2}}{2\sigma_{\text{ran}}^{2}}\right]$ (2.3)
This function models the absolute log-nomral distribution of redshift catastrophics ($\Delta\hat{v}_{\rm{catas}}=\log_{10}(|\Delta v_{\rm{catas}}|)$) with parameters $\mu_{\text{ran}}$, $\sigma_{\text{ran}}^{2}$ and $v_{0}$ to characterize. In our study, we focus on the random pattern while varying the contamination rate $f_{c}$ to generalize catastrophics patterns for different surveys.
2.3 Combined Errors for Slitless Redshift Surveys
Space missions such as Euclid [13], Roman [38] and CSST (scheduled for launch in 2027) [39] employ slitless spectroscopy to map the 3D distribution of galaxies over wide areas of sky. Early datasets from Euclid (Quick Data Release 1; [34]) are now publicly available, providing initial evaluations of the performance of the redshift measurement. Unlike traditional slit-based spectroscopy that captures individual targets using fibers, slitless spectroscopy disperses the light from all sources in the field simultaneously, enabling efficient and continuous spectroscopic observations without the need for target pre-selection [29, 17]. This approach allows for high sky coverage and dense sampling of galaxies in the redshift range relevant to LSS studies.
However, the slitless design magnifies the redshift errors. Euclid spectroscopic redshifts are derived from spectra provided by the Near Infrared Spectro-Photometer (NISP), which operates with a resolution of $R\simeq 500$ [29]. In comparison, the ground-based DESI survey uses instruments with resolution $R\sim 2000$ to $5500$ [10]. The lower spectral resolution in slitless survey leads to broader spectral lines and consequently increased redshift uncertainty. Additionally, the lack of spatial isolation between sources causes spectral contamination, typically through cross-contamination and self-contamination [17]. Cross-contamination arises from overlapping spectrograms of neighboring sources (redshift interlopers; [23]), while self-contamination results from the degeneracy between spatial and spectral information along the dispersion direction. These combined effects lead to a higher rate of redshift catastrophics in the sample. To address these challenges, we adopt a hypothetical redshift error that combines both redshift uncertainty and catastrophics to simulate the expected observational systematics in slitless surveys. The model assumes larger uncertainty than those typical of ELGs (i.e., LRG-like) and a higher rate of catastrophic errors (5% with $|\Delta v|>1000\,\rm{km}\cdot s^{-1}$). This hypothetical redshift error will help us to assess their influences on cosmological analyses obtained from space slitless surveys.
3 Contaminated Mocks
Mocks label clean dv-LRG dv-QSO-G dv-QSO-L dv-ELGcatas dv-slitless
$\sigma_{\Delta v}$$w_{\Delta v}$ ($\rm{km}\cdot s^{-1}$) - 85.7 200 92.7 - 85.7
$f_{c}$ - - - - 1% 5%
Table 1: Summary of contaminations and mock catalogs used in our study. Mocks include uncontaminated (clean) one and those affected by redshift errors. Redshift uncertainty is modeled using Gaussian (G) or Lorentzian (L) profiles, characterized by $\sigma_{\Delta v}$ or $w_{\Delta v}$ respectively. We consider uncertainties for LRG-like Gaussian (dv-LRG), QSO-like Gaussian (dv-QSO-G) and QSO-like Lorentzian (dv-QSO-L), using the magnitudes reported from DESI EDR repeat observations [26]. Redshift catastrophic failures are based on analyses of DESI DR1 ELG repeat observations [21], with $f_{c}$ being the contamination rate. We consider $f_{c}=1\%$ with ELG-like catastrophic pattern (dv-ELGcatas) in DESI survey. For slitless-like errors (dv-slitless), we assume combined errors with LRG-like uncertainty and $f_{c}=5\%$ ELG-like catastrophics to mimic the expected conditions in Euclid [17]. The corresponding redshift error distributions for each contaminated mock are shown in Figure˜1.
Figure 1: Comparisons of redshift errors distributions. Left: Typical LRG-like and QSO-like redshift uncertainties, modeled using Gaussian (G) or Lorentzian (L). Right: ELG-like catastrophics in DESI survey and hypothetical slitless-like errors for Euclid, which combines LRG-like redshift uncertainty (left part) and $f_{c}=5\%$ ELG-like catastrophic contamination (right part). These redshift error distributions are used to generate the additional velocity dispersion and construcut our contaminated mock catalogs.
We use the Quijote $N$-body simulation suite to construct our mock catalogs [40]. The Quijote fiducial simulations assume a flat $\Lambda\rm{CDM}$ universe, with cosmology consistent with the Planck 2018 result [41]: $h=0.6711,\,\Omega_{\mathrm{m}}=0.3175,\,\Omega_{\mathrm{b}}=0.049,\,\sigma_{8}=0.834,\,n_{s}=0.9624,\,\sum m_{\nu}=0.0\,{\rm{eV}},\,w=-1$. To study the impact of redshift errors on neutrino mass measurement, we also include the ‘Mnu_p’ catalog with $\sum m_{\nu}=0.1,\mathrm{eV}$. In Quijote simulation, each realization contains $512^{3}$ CDM particles with mass $m_{p}\approx 6.5\times 10^{11}\,M_{\odot}$, evolved within a comoving box of volume $1\,h^{-3}\rm{Gpc}^{3}$. To match the redshift range of interest, we use the snapshots at redshift $z=1.0$. A mass selection with $13.1<\log_{10}(M_{h}/(h^{-1}M_{\odot}))<13.5$ is applied to ensure that the truncated halos represent the dominant structures contributing to the galaxy clustering. We use in total 500 realizations to ensure a minimal statistical noise and allow us to robustly assess the potential biases introduced by redshift errors.
The halo catalogs include information such as the Cartesian positions $X,Y,Z$, the corresponding peculiar velocities $\bm{v}_{\rm{pec}}$, and the halo masses, $M_{h}$. Assuming the line-of-sight is along the $x$-direction, we use the $x$-component of the peculiar velocity, $v_{{\rm{pec}},x}$, to transform halo positions from real space to redshift space by $X_{\rm{rsd}}=X_{\rm{real}}+v_{{\rm{pec}},x}\,(1+z)/H(z)$. The clean mock catalogs are built using the peculiar velocity without modification. To build contaminated mocks with redshift errors, we add an additional velocity shift $\Delta v_{\rm{error}}$ to $v_{{\rm{pec}},x}$. This $\Delta v_{\rm{error}}$ simulates the line-of-sight velocity dispersion introduced by redshift errors.
To generate the $\Delta v_{\rm{error}}$ distribution, we use different profiles to simulate redshift errors for different surveys. For redshift uncertainty, we use the magnitudes reported from repeat observations in DESI Early Data Release (EDR) [26]. Specifically, we apply Gaussian profiles (defined in Eq.˜2.1) with $\sigma_{\Delta v}=85.7\,\rm{km}\cdot s^{-1}$ for LRG-like smearing and $\sigma_{\Delta v}=200\,\rm{km}\cdot s^{-1}$ for QSO-like smearing. To simulate the extended tails observed in QSO redshift errors, we also consider a Lorentzian profile (defined in Eq.˜2.2), with $w_{\Delta v}=95.7,\rm{km}\cdot s^{-1}$, which is truncated with $\Delta v_{\rm{smear}}<2000\,\rm{km}\cdot s^{-1}$ to avoid divergence. A summary of the configurations used for the contaminated mocks is in Table˜1, with the comparison of distributions shown in the left panel of Figure˜1. In all cases, we assume no systematic redshift offset (i.e., $\mu=0$ and $p=0$), consistent with observational findings that show minimal bias in the mean redshift determination. The redshift smearing effects for ELG are relatively small and therefore not included in this redshift uncertainty study [21].
For redshift catastrophics, we adopt a model based on DESI DR1 ELG repeat observations [21]. The velocity shifts caused by catastrophics follow a symmetric distribution centered at zero, with each half modeled by the profile $\mathcal{LN}(\Delta\hat{v},|,0.64,0.25^{2},6.62)$ defined in Eq.˜2.3. We set $f_{c}=1\%$ to represent a ELG-like catastrophics, corresponding to the upper limit of contamination rate in eBOSS and DESI galaxies. As discussed in Section˜2.3, to mimic redshift errors expected in slitless surveys, we consider a more pessimistic scenario combining LRG-like Gaussian smearing ($\sigma_{\Delta v}=85.7\,\rm{km}\cdot s^{-1}$) with $f_{c}=5\%$ ELG-like catastrophics. The contaminated mocks are summarized in Table˜1 and the distributions of ELG-like catastrophics and sliltless-like errors are shown in the right panel of Figure˜1.
4 Methods
4.1 Power Spectrum Estimator
We estimate the power spectrum of mock catalogs using the weighted galaxy density field with a convenient normalization [42]:
$F(\textbf{r})=\frac{w(\textbf{r})\left[n_{g}(\textbf{r})-\alpha n_{s}(\textbf{r})\right]}{A^{1/2}}$ (4.1)
where $w(\textbf{r})$ is the weight function. $n_{g}$ and $n_{s}$ are the observed number densities of galaxy and random catalogs, respectively. The factor $\alpha=N_{g}/N_{s}$ normalizes the contribution from the random, and the normalization factor $A$ is defined with the expected mean space density of galaxies $\bar{n}(\textbf{r})$ as $A=\int d\textbf{r}\left[\bar{n}(\textbf{r})w(\textbf{r})\right]^{2}$. The estimator of power spectrum multipoles $P_{\ell}$ is given by [42, 43]:
$P_{\ell}(\vec{k})=(2\ell+1)\int\frac{d\Omega_{k}}{4\pi}\left[\int d\textbf{r}_{1}\int d\textbf{r}_{2}F(\textbf{r}_{1})F(\textbf{r}_{2})\times e^{i\vec{k}\cdot(\textbf{r}_{1}-\textbf{r}_{2})}\mathcal{L}_{\ell}(\hat{\textbf{k}}\cdot\hat{\textbf{r}}_{h})-\rm{SN}_{\ell}\right]$ (4.2)
[…]
4.2 Galaxy Clustering Models
4.2.1 Effective-field theory of LSS
Effective field theory (EFT) has become a powerful tool to model the LSS of the Universe. By studying small-scale nonlinear physics, the EFT of LSS extend the standard perturbation theory to scales beyond the breakdown of the fluid and single-stream approximations [45, 46, 47]. With loop corrections and improved modeling of biased tracers with redshift-space distortions, EFT enables reliable predictions of the galaxy power spectrum multipoles up to $k\approx 0.25\,h\,\rm{Mpc}^{-1}$ in redshift space [48, 49, 50, 51]. In general, the galaxy redshift-space power spectrum in EFT framework is expressed as [8]:
[…]
4.2.2 Full-Shape Fitting Approaches
We adopt two fitting strategies to perform full-shape cosmological inference: ShapeFit compression [56] and Full-Modeling. These approaches have been extensively tested and validated in analyses of the BOSS and DESI surveys [57, 58, 59, 60, 61].
Compression analysis focuses on fitting a set of compressed physical parameters that describe the shape and the anisotropies of the galaxy power spectrum. These include the $f\sigma_{8}$ and the scaling parameters $\alpha_{\parallel}$ and $\alpha_{\perp}$, which are often reparameterized into the isotropic dilation parameter $\alpha_{\mathrm{iso}}\equiv\alpha_{\parallel}^{1/3}\alpha_{\perp}^{2/3}$ and the Alcock-Paczynski (AP) parameter $\alpha_{\mathrm{ap}}\equiv\alpha_{\parallel}\alpha_{\perp}^{-1}$. ShapeFit further introduces two shape-related parameters, $m$ and $n$, to capture information from the early-time transfer function. In this work, we vary only one shape parameter $m$, while keeping $n$ fixed, and we report ShapeFit results relative to a reference cosmology, specifically using $dm\equiv m-1$ and $df\equiv f/f^{\rm{fid}}$.
Another fitting approach, known as Full-Modeling analysis, avoids the compression steps and directly fits cosmological models to the power spectrum signals. In this way, the parameter dependence of the linear power spectrum along and geometrical information are not kept separate. This allows for a direct comparison between theoretical predictions and observational data, leading direct constraints on cosmological parameters. In our baseline fitting procedure, we sample five cosmological parameters: $h$, $\omega_{\rm cdm}$, $\omega_{b}$, $\ln(10^{10}A_{s})$, $n_{s}$. These parameters are directly used in theoretical models, which facilitates the calculation and a direct comparison of cosmological biases introduced by redshift errors.
In the context of redshift error studies, ShapeFit provides a direct view of how redshift errors distort the shape of observed power spectrum, as reflected by their impact on compressed physical parameters. This makes the interpretation of redshift errors effect straightforward. The Full-Modeling approach, on the other hand, gives cosmological constraints, allowing us to quantify the biases introduced by redshift errors in final cosmological parameter constraints.
4.3 Clustering effect of redshift errors
4.3.1 Redshift uncertainty damping
As pointed out in the eBOSS QSO analysis [62, 24, 20], a straightforward approach to model the impact of redshift uncertainty is to introduce an additional Gaussian damping term:
$P_{g,\,s}(k,\mu)=P_{\rm{novir}}(k,\mu)\,F_{\rm{FOG}}(k,\mu)\exp\left[-(k\mu\sigma^{\prime})^{2}\right]$ (4.4)
Here, $P_{\rm{novir}}$ is the non-linear kaiser power spectrum and $F_{\rm{FOG}}$ is finger-of-god (FOG) factor. In the last Gaussian damping term, the parameter $\sigma^{\prime}$ quantifies the scale of redshift uncertainty and is related to the equivalent velocity smearing with $\sigma^{\prime}=\Delta v_{\rm{smear}}*{(1+z)}/{H(z)}$ [62]. Using a Taylor expansion of the Gaussian exponential term:
$\exp\left[-(k\mu\sigma^{\prime})^{2}\right]\,\rightarrow 1+a_{2}(k\mu)^{2}+a_{4}(k\mu)^{4}+\ldots$ (4.5)
We note that the second and third terms scale as $P(k)(k\mu)^{2}$ and $P(k)(k\mu)^{4}$, respectively. These terms show similar structures to those of EFT counterterms in Eq.˜4.3. In the EFT of LSS, the general form of such counterterms can be derived from a moment expansion of the redshift-space power spectrum [51], as well as from comparisons with phenomenological models [6]. These studies suggest that this part of EFT counterterms, parameterized by $\alpha_{2}$ and $\alpha_{4}$ particularly, are capable of absorbing nonlinear effects introduced by redshift uncertainty. We will demonstrate this explicitly in Section˜5 with detailed explanation in Appendix˜B.
4.3.2 Redshift catastrophics correction
Redshift catastrophics have been addressed in theoretical models for redshift surveys like WiggleZ [14] and recent Euclid [23]. Assuming that the redshift catastrophics contamination rate is $f_{c}$, an approximated correction is to multiply the EFT galaxy power spectrum by a factor of $(1-f_{c})^{2}$:
$P^{\texttt{EFT}+f_{c}}_{g,\,s}(k,\mu)\approx(1-f_{c})^{2}P^{\texttt{EFT}}_{g,\,s}(k,\mu)$ (4.6)
The validity of this approximation depends on both $f_{c}$ and the typical displacement caused by catastrophics. Based on our mock catalogs study with contamination $f_{c}=1\%$ and $f_{c}=5\%$, we find that the relative difference between the correction factor, $R_{\rm{corr}}=(1-f_{c})^{2}$, and the ratio from the mocks, $R_{\rm{mock}}$, satisfies:
$\Bigg{\lvert}\frac{R_{\rm{corr}}(k)-R_{\rm{mock}}(k)}{R_{\rm{mock}}(k)}\Bigg{\lvert}<10\%,\quad\rm{For}\ k\gtrsim 0.02\,h\,\rm{Mpc}^{-1}$ (4.7)
This suggests that the approximation is sufficiently accurate for full-shape power spectrum analyzes, especially considering that additional nuisance parameters in the EFT can absorb residual mismatches. However, this correction is not applied in the standard EFT full-shape analysis. This is because, as we will demonstrate in Section˜5.2.2, introducing this free parameter $f_{c}$ in the fitting process will lead to the loss of constraining power on certain cosmological parameters.
4.4 Cosmological Inference and Covariance Matrices
For cosmological inference, Bayesian statistics is employed to infer constraints on physical or cosmological parameters. Assuming the evidence from the data is normalized to one, the posterior probability is expressed as $P(\Theta|\mathcal{D})\propto\mathcal{L}(\mathcal{D}|\Theta)P(\Theta)$, with $\Theta$ being the set of input parameters, including compressed parameters $\left\{\alpha_{\mathrm{iso}},\;\alpha_{\mathrm{ap}},\;df,\;dm\right\}$ by ShapeFit or cosmological parameters $\left\{h,\;\omega_{cdm},\;\omega_{b},\;\rm{ln}(10^{10}A_{s}),\;n_{s}\right\}$ by Full-Modeling. It also contains nuisance parameters considered in the EFT model: $\left\{b_{1},\;b_{2},\;b_{s},\;\alpha_{0},\;\alpha_{2},\;\alpha_{4},\;\rm{SN}_{0},\;\rm{SN}_{2}\right\}$. Assuming that we have Gaussian-distributed data, the likelihood function can be expressed as $\mathcal{L}(\Theta)\propto e^{-\chi^{2}(\Theta)/2}$. The $\chi^{2}$ statistic quantifies the goodness of fitting and is computed based on the difference between the observed data and the model predictions, which reads as:
$\chi^{2}(\Theta)=(\mathcal{D}_{\rm{data}}-\mathcal{D}_{\rm{model}}(\Theta))^{T}\,\mathcal{C}^{-1}\,(\mathcal{D}_{\rm{data}}-\mathcal{D}_{\rm{model}}(\Theta))$ (4.8)
In our study, the vector $\mathcal{D}$ includes a combination of the power spectrum monopole $P_{0}(k)$ and quadrupole $P_{2}(k)$. The subscript ’data’ and ’model’ refers to the measurements from observations (here mocks) and theoretical model predictions. The covariance matrix $\mathcal{C}$ is extracted from Quijote 500 mock catalogs. It incorporates the variances observed in the data and is normalized to the number of the datasets, $N_{d}$, with the formulation:
$\mathcal{C}_{i,j}=\frac{1}{N_{d}-1}\ \,\langle(\mathcal{D}_{i}-\overline{\mathcal{D}}_{i})(\mathcal{D}_{j}-\overline{\mathcal{D}}_{j})\rangle$ (4.9)
Where $\mathcal{D}_{i}$ signifies the power spectrum from a single realization and $\overline{\mathcal{D}}_{i}$ is ensemble mean. In Quijote suite, the volume of each realization is $V_{1}=1\,h^{-3}\rm{Gpc}^{3}$. To model the reduced cosmic variance expected from galaxy surveys, we use the rescaled covariance by volume as: $\mathcal{C^{\prime}}=({V_{1}}/{V_{n}})\,\mathcal{C}$, where $V_{n}$ is the target survey volume. In our analyses, we rescale the covariance to an effective volume of $V_{25}=25\,h^{-3}\rm{Gpc}^{3}$ to match the expected survey volumes of DESI and Euclid.
5 Results
5.1 Impact of redshift errors on clustering statistics
Figure 2: Comparison of power spectrum monopole (left) and quadrupole (right) between clean mocks and mocks contaminated by redshift uncertainty. The contaminated mocks are generated to mimic the observational effect of LRG-like and QSO-like smearing. The error bars, $\sigma$, are rescaled to match an effective survey volume of $V_{25}$. The middle subpanels display the power spectrum ratios, $R_{\ell}(k)$, between the contaminated and clean mocks. The lower subpanels show the their residuals in units of $\sigma$, with the dotted line indicating the $2\sigma$ threshold. We see significant clustering deviations ($2\sigma$) due to QSO-like Gaussian or Lorentzian smearing for $k\gtrsim 0.10\,h\,\rm{Mpc}^{-1}$.
Figure 3: Same plot as Figure˜2 comparing clean and mocks contaminated by ELG-like catastrophics and slitless-like errors. The middle subpanels show that the power spectrum ratios $R(k)$ and expected reduction $(1-f_{c})^{2}$ due to the redshift catastrophics (dash line). Slitless-like errors introduce an overall suppression of power spectrum amplitude by $10\%{-}12\%$, exceeding $3\sigma$ level in both monopole and quarupole.
The impact of redshift errors on the clustering statistics is presented in Figure˜2 and Figure˜3. For redshift uncertainty, it introduces a diffusive effect that systematically suppresses the power spectrum amplitude, with the suppression becoming stronger at smaller scales (i.e., larger $k$). As discussed in Section˜4.3.1, this effect can be modeled by a Gaussian damping term. In the middle subpanels of Figure˜2, we plot the ratios of power spectra, $R(k)$, between the contaminated mocks and clean mocks and we find that $R(k)$ exhibits Gaussian-like profiles, with larger $\Delta v_{\rm{smear}}$ resulting in stronger damping of the power spectrum. For an effective survey volume of $V_{25}$, the LRG-like smearing leads to minimal impacts on the clustering, with power spectrum difference $\Delta P\lesssim 2\sigma$ (see lower subpanels in Figure˜2). For QSO-like redshift uncertainty, the Lorentzian and Gaussian smearing introduce a comparable suppression, with $\Delta P>2\sigma$ for $k\gtrsim 0.10\,h\,\rm{Mpc}^{-1}$ in both the monopole and quadrupole.
Redshift catastrophics affect only a small fraction of objects but cause significant shifts of their positions along the line-of-sight. In Section˜4.3.2, we show that this distortion leads to an overall reduction on the power spectrum, scaling approximately as $(1-f_{c})^{2}$. As illustrated in Figure˜3, the ELG-like catastrophics with $f_{c}=1\%$ reduce the clustering signal to $\sim 98\%$ of its original amplitude, with deviations below $3\sigma$. For slitless-like errors (i.e., $\sigma_{v}=85.7\,\rm{km}\cdot s^{-1}$ and $f_{c}=5\%$), the power spectrum amplitude from contamianted mocks is reduced to $(1-5\%)^{2}=90.25\%$, i.e., a suppression of $10\%$. An additional deviation from the $(1-f_{c})^{2}$ arises due to the smearing effect introduced by redshift uncertainty. In this case, the overall suppression of slitless-like errors is more significant, exceeding $3\sigma$ for $k\gtrsim 0.02\,h\,\rm{Mpc}^{-1}$ in the monopole and $k\gtrsim 0.10\,h\,\rm{Mpc}^{-1}$ in the quadrupole. Such deviations, if uncorrected, can bias the cosmological inference. In the following subsections, we describe their influences when running the inference in different cosmological cases: $\Lambda\rm{CDM}$, $w_{0}w_{a}\rm{CDM}$ and $\nu\Lambda\rm{CDM}$.
5.2 Baseline Fitting
Our baseline full-shape analysis fits the power spectrum monopole and quadrupole over the range $0.02\leq k\leq 0.2\,h\,\rm{Mpc}^{-1}$, using the standard EFT model without applying the correction factor of redshift catastrophics. We adopt two fitting strategies, ShapeFit and Full-Modeling, to extract cosmological information. The results are presented within the standard flat-$\Lambda\rm{CDM}$ model, assuming a BBN prior on baryon abundance $\Omega_{\mathrm{b}}$, and wide Gaussian prior on spectral index $n_{s}$. The cosmological inference is performed using the desilike [^3] https://github.com/cosmodesi/desilike package and parameter priors are set to the same as the those in DESI DR1 full-shape analysis [61]. The parameter space is sampled with the code emcee [63]. All plots and confidence intervals are generated using the GetDist  package [64]. Beside the baseline full-shape analysis, we also evaluate the impact of redshift errors on BAO fitting. These results are discussed in Appendix˜A.
5.2.1 Cosmological influence
Figure 4: Distributions of the best-fit values for ShapeFit physical parameters. Scatter plots compare results from clean and contaminated mocks in Table˜1, with each color representing a different error type. The corresponding marginalized histograms are shown in the right and top subpanels. Each point represents one fit to the mean of 25 random mocks, with the covariance rescaled to volume $V_{25}$. Overall, the parameter estimates remain consistent across most cases, except for $df$, which shows a notable deviation in the mocks contaminated by slitless-like errors.
Figure 5: Same plots as Figure˜4 for Full-Modeling cosmological parameters. As before, results from clean and contaminated mocks in Table˜1 are compared. Only $\ln(10^{10}A_{s})$ shows a notable deviation in mocks contaminated by slitless-like errors.
In this section, we assess the impact of redshift errors on cosmological inference. These results are based on 200 independent fits to the Quijote mocks catalogs in Table˜1. For each fit, we use the mean power spectrum derived from 25 randomly selected mocks, with a covariance matrix rescaled to the volume $V_{25}$. Comparisons of the best-fit parameters between analyses from clean and contaminated mocks are presented in Figure˜4 for ShapeFit and in Figure˜5 for Full-Modeling. The corresponding $\chi^{2}$ comparison, provided in Appendix˜B, confirms that all our fits achieved comparable goodness-of-fit.
In Figure˜4 and Figure˜5, the upper panels illustrate the impact of redshift uncertainty on cosmological inference. The fitting results of ShapeFit and Full-Modelling indicate that there is no significant shift in parameter estimates (<5%). As discussed in Section˜4.3.1, the suppression of the power spectrum caused by redshift uncertainty is expected to be absorbed by the EFT counterterms. In Appendix˜B, we plot the posterior distributions for all nuissance parameters from a representative fit, which shows negative shifts in counterterm parameters $\alpha_{2}$ and $\alpha_{4}$ for contaminated case. This is consistent with our theory explanations. Additionally, by comparing QSO-like uncertainties modeled with Gaussian and Lorentizan profile, we find that, as long as they have the same magnitude of smearing, the presence of extended tail does not significantly affect cosmological measurements in full-shape analysis.
The lower panels of Figure˜4 and Figure˜5 assess the impact of redshift catastrophics and hypothetical slitless-like errors on cosmological inference. The ELG-like catastrophics with $f_{c}=1\%$ show negligible bias in parameter estimates (with the largest discrepancy $6\%$), consistent with the previous findings in [21]. In contrast, slitless-like errors with $f_{c}=5\%$ show noticeable biases, particularly in estimating the fractional growth rate parameter $df$ in ShapeFit and the log primordial amplitude parameter $\ln(10^{10}A_{s})$ in Full-Modeling. Both parameters are systematically underestimated by $6\%{-}16\%$ compared to clean case. We identify that this negative shift arises from the $5\%$ catastrophics contamination in slitless-like errors, which significantly suppresses the power spectrum amplitude (see Figure˜2). If not adequately modeled in theory, this effect could bias the estimates of amplitude-related parameters. Nevertheless, as discussed in Section˜4.3.2, a simple correction by multiplying the power spectrum by a factor of $(1-f_{c})^{2}$ can mitigate this suppression. In the following section, we incorporate this correction into the cosmological analysis to further investigate the influence of redshift catastrophics.
5.2.2 Catastrophics-correction for slitless survey
In this test, we present Full-Modeling fitting results based on the mean power spectrum calculated from all 500 mocks, with the covariance matrix rescaled to volume $V_{25}$. By averaging over 500 mocks, we minimize the impact of cosmic variance, so any observed deviations in parameter estimates can be more confidently attributed to redshift errors. The left panel of Figure˜6 shows the posterior distributions of cosmological parameters using the baseline EFT fit for three contamination cases: clean (no contaminations), ELG-like catastrophics ($f_{c}=1\%$), and slitless-like errors (i.e., $\sigma_{v}=85.7\rm{km}\cdot s^{-1}$ and $f_{c}=5\%$). The corresponding mean values with 68% confidence intervals for $\ln(10^{10}A_{s})$ are summarized in Table˜2. Compared the $\ln(10^{10}A_{s})$ constraint from clean case, we find a small shift of $0.36\sigma$ by ELG-like catastrophics and a significant deviation of $2.20\sigma$ by slitless-like errors. These results are consistent with our earlier findings in Section˜5.2.1.
To incorporate the effect of catastrophics in the clustering model, we extend the baseline EFT model by adding an additional parameter $f_{c}$ and multiplying the factor $(1-f_{c})^{2}$ to the galaxy power spectrum as described in Eq.˜4.6. We refer to the $\texttt{EFT}+{\rm{free}}\,f_{c}$ model, when the parameter $f_{c}$ varies during sampling with a flat prior $f_{c}\in[0,1]$. The posterior distributions, including the variation of $f_{c}$, are shown in the right panel of Figure˜6 (orange), with the corresponding constraint on $\ln(10^{10}A_{s})$ reported in Table˜2. The $\texttt{EFT}+{\rm{free}}\,f_{c}$ model is able to recover an unbiased mean value of $\ln(10^{10}A_{s})$ for slitless-like errors, with a discrepancy below $0.1\sigma$ compared to the clean case. However, this improvement comes at the cost of loss in constraining power: we measure a $\sim 60\%$ degradation in the $\ln(10^{10}A_{s})$ constraint. This loss is due to the strong degeneracy between $f_{c}$ and $\ln(10^{10}A_{s})$, as illustrated in the $\ln(10^{10}A_{s}){-}f_{c}$ plane shown in the right panel of Figure˜6. When fixing the parameter $f_{c}$ to its expected value, referred to the $\texttt{EFT}+{\rm{fixed}}\,f_{c}$ model (yellow), we successfully recover a constraining power comparable to that of the clean case. The estimated mean value of $\ln(10^{10}A_{s})$ inferred by the $\texttt{EFT}+{\rm{fixed}}\,f_{c}$ model shows a slightly larger deviation ($\sim 0.85\sigma$), likely due to the combined effects of redshift uncertainty in the slitless-like errors and the approximation applied in the correction model (see the $R(k)$ comparison in Figure˜3).
This test is informative for slitless spectroscopy survey like Euclid, which aim to constrain cosmological parameters via galaxy power spectra. As we have shown that this hypothetical redshift errors can affect our cosmological inference, an accurate estimation of the catastrophics rate $f_{c}$ and the implementation of the catastrophics-corrected model, either with $f_{c}$ fixed or varied under a narrow Gaussian prior, will be essential to ensure unbiased constraints.
Figure 6: Posterior distributions for Full-Modeling cosmological parameters. Left panel: Constraints using the baseline EFT model to fit clean mocks (dark blue), mocks contaminated by ELG-like catastrophics ($f_{c}=1\%$; purple), and slitless-like errors ($\sigma_{v}=85.7\rm{km}\cdot s^{-1}$ with $f_{c}=5\%$; red). Right panel: Comparison between clean results with slitless-like errors using the extended EFT model when freeing $f_{c}$ ($\texttt{EFT}+{\rm{free}}\,f_{c}$; orange) and with $f_{c}$ fixed to its expected value ($\texttt{EFT}+{\rm{fixed}}\,f_{c}$; yellow). Contours show 68% and 95% confidence intervals with black dash lines indicating the true values. Results are derived from fits to the mean power spectrum monopole and quadrupole of 500 mocks using the covariance matrix rescaled to $V_{25}$. The most notable change is in $\ln(10^{10}A_{s})$ estimate, for which the constraints are summarized in Table˜2.
Theory EFT baseline $\texttt{EFT}+{\rm{free}}\,f_{c}$ $\texttt{EFT}+{\rm{fixed}}\,f_{c}$ Mocks clean dv-ELGcatas dv-slitless dv-slitless dv-slitless $\ln(10^{10}A_{s})$ $3.125\pm 0.084$ $3.095\pm 0.081$ $2.941\pm 0.085$ $3.13\pm 0.14$ $3.054\pm 0.080$
Table 2: Constraints on $\ln(10^{10}A_{s})$ for clean and contaminated mocks analyzed by different theoretical models. Table shows the mean value with 68% confidence intervals. In EFT baseline fit, the ELG-like catastrophics (dv-ELGcatas) and sliltless-like errors (dv-slitless) show deviations of $0.36\sigma$ and $2.20\sigma$, respectively, compared to the clean case. By multiplying the factor $(1-f_{c})^{2}$ to the galaxy power spectrum, the extended model $f_{c}$, $\texttt{EFT}+{\rm{free}}\,f_{c}$ recovers an unbiased value of $\ln(10^{10}A_{s})$ but increases its error bar by 60% due to parameter degeneracies. In contract, the $\texttt{EFT}+{\rm{fixed}}\,f_{c}$ model restore the posterior without the loss of constraining power.
5.3 Beyond $\Lambda\rm{CDM}$ cosmologies
Recent spectroscopic surveys such as DESI have extended full-shape power spectrum analyses beyond the standard $\Lambda\rm{CDM}$ framework to test cosmological models with evolving dark energy ($w_{0}w_{a}\rm{CDM}$) and massive neutrinos ($\nu\Lambda\rm{CDM}$). Assuming a flat Universe and combining the full-shape and BAO analysis with external datasets, the DESI DR2 analysis finds a mild preference for a dynamical dark energy model with $w_{0}>-1$ and $w_{a}<0$  [5] and places a tight constraint on the total neutrino mass, $\sum m_{\nu}<0.0642\,\rm{eV}\ (95\%CL)$ [65]. This section investigates how redshift errors affect the measurements of dark energy equation of state $w_{0}$-$w_{a}$ and $\sum m_{\nu}$ through the full-shape analysis. In the following, our results show fits to the mean power spectrum of 500 mocks with the covariance rescaled to volume $V_{25}$.
5.3.1 Evolving dark energy
Figure 7: Parameter constraints on $w_{0}w_{a}\rm{CDM}$ model for clean and contaminated mocks. The horizontal error bars indicate mean and the 68% credible levels of the corresponding posterior. The shaded region show the $1\sigma$ constraints from the clean mocks. The open circles represent the MAP values. While redshift errors do not introduce bias in the measurement of dark energy parameters $w_{0}$, $w_{a}$, they potentially lead to stronger projection effects (deviation between MAP and mean of the posterior), particularly in cases with larger redshift uncertainty.
We assess the influence of redshift errors on the measurement of evolving dark energy using the Chevallier-Polarski-Linder (CPL) parametrization $w(a)=w_{0}+w_{a}(1-a)$ [66]. The cosmological fits are performed using the Quijote fiducial mocks described in Section˜3, which assume a $\Lambda\rm{CDM}$ model with $w_{0}=-1$ and $w_{a}=0$. The resulting parameter constraints are shown as the horizontal lines in Figure˜7. From the comparisons, we find that redshift errors do not introduce significant bias in the measurement of dark energy properties. However, we know that using full-shape analyses alone to constraint $w_{0}w_{a}\rm{CDM}$ model suffer from projection effects [9]. To study this aspect, we use iminuit [67] to estimate the maximum a posteriori (MAP) value, starting from the maximum log-posterior points found in MCMC chains. In Figure˜7, we find offsets between the mean of the marginalised posteriors (solid circles) and MAP values (open circles) for parameters $h$, $w_{0}$ and $w_{a}$, with the offsets being most pronounced for mocks contaminated by QSO-like Lorentzian smearing. Since the redshift errors impact the clustering pattern similarly with the behavior of certain parameters (cosmological or nuissance), their presence could complicate the parameter degeneracies and thus amplify projection effects. Such impact should be investigated in more detail for redshift surveys aiming to constrain dark energy by full-shape analysis.
5.3.2 Massive neutrinos
Figure 8: 1D marginalized posterior distributions of sum of neutrino mass, $\sum m_{\nu}$, for uncontaminated mock (solid darkblue line) and contaminated (dashed lines) mock. Left panel show cases with LRG-like (purple), QSO-like (Gaussian, light purple; Lorentzian, orange) redshift uncertainties. Right panel use contaminated mocks with ELG-like catastrophic ($f_{c}=1\%$, blue) and slitless-like errors ($\Delta v_{\rm{smear}}=85.7\rm{km}\cdot s^{-1}$ with $f_{c}=5\%$, green). The shaded region indicates the 95% confidence level from the clean mocks, while the vertical dotted line marks the true value $\sum m_{\nu}=0.1\,\rm{eV}$.
Catalogs clean dv-LRG dv-QSO-G dv-QSO-L dv-ELGcatas dv-slitless $\sum m_{\nu}$ $0.80\,\rm{eV} <$$1.07\,\rm{eV} <$$1.08\,\rm{eV} <$$1.03\,\rm{eV} <$$0.88\,\rm{eV} <$$0.83\,\rm{eV}$
Table 3: Comparison of 95% confidence upper bounds on the sum of neutrino masses, $\sum m_{\nu}$. Results are obtained using Full-Modeling analyses with $\sum m_{\nu}$ varied and fitting power spectrum monopole and quadrupole of 500 mocks with covariance rescaled to $V_{25}$. Both clean and contaminated mock catalogs are considered with $\sum m_{\nu}=0.1,\mathrm{eV}$. We find that redshift uncertainties degrade the neutrino mass constraint by $\sim 30\%$, whereas ELG-like redshift catastrophics and slitless-like errors show negligible effect.
For the $\nu\Lambda\rm{CDM}$ case, we use the Quijote Mnu_pMnu\_p catalogs with $\sum m_{\nu}=0.1\,\rm{eV}$ to build our clean and contaminated mock catalogs. During the fitting, we allow $\sum m_{\nu}$ to vary with a flat prior of $\mathcal{U}(0.0,5.0)\,\rm{eV}$, while keeping other settings the same as the baseline. The marginalized posteriors of $\sum m_{\nu}$ for all mocks analyses are displayed in Figure˜8 ana the upper limits of $\sum m_{\nu}$ at 95% confidence level are presented in Table˜3. We first note that using full-shape analysis alone to constrain the $\sum m_{\nu}$ has slow convergence and large uncertainties, particularly for the low neutrino mass case. Nevertheless, we find that redshift uncertainty introduces extended tails in the $\sum m_{\nu}$ posterior distribution, leading to a weaker constraint on the neutrino mass. Compared to the clean case ($0.80\,\rm{eV}$), the 95% upper limits increase by $\sim 28$ to 35% when redshift uncertainty is included. This degradation occurs because both massive neutrinos and redshift uncertainty have similar suppression on small-scale power spectrum. Massive neutrinos do so through their thermal velocities and free-streaming, while redshift uncertainty smear the clustering structure along the line of sight. This degeneracy complicates the convergence and leads to weaker constraints on neutrino mass. In contrast, we find no noticeable bias in the $\sum m_{\nu}$ measurements due to ELG-like catastrophic failures or slitles-like redshift errors.
6 Conclusions
In this paper, we investigate the impact of spectroscopic redshift errors on the cosmological measurements. Section˜2 introduces two main types of redshift errors, redshift uncertainty and redshift catastrophic failures, which can affect redshift measurements in both ground-based (fiber-fed) and space-based (sltiless) spectroscopic surveys such as eBOSS, DESI and Euclid [35, 21, 17]. While redshift uncertainty refers to small errors present in all redshift measurements, redshift catastrophic failures correspond to large deviations from the true values in a small fraction of the measurements. Both types of error distort the observed clustering pattern and can lead to potential biases in the inferred cosmological parameters.
The effect of redshift errors can be quantified and modeled using repeat redshift observations. Based on this, in Section˜3 we use the Quijote simulation suite to constructed our contaminated mock catalogs. For redshift uncertainty, we consider three cases: LRG-like Gaussian, QSO-like Gaussian and QSO-like Lorentzian smearings, with the velocity dispersions informed by DESI EDR studies [26]. For redshift catastrophics, we use a log-normal Gaussian distribution [21] with contamiantion rate $f_{c}=1\%$ to model ELG-like catastrophics and $f_{c}=5\%$ for hypothetical slitless-like errors. The slitless-like errors also combines LRG-like Gaussian smearing to simulate the observing condition expected in slitless spectroscopic survey such as Euclid. The following Section˜4 describes the methodologies to perform the full-shape power spectrum analysis. We use clustering models based on the effective field theory (EFT) of LSS and implement both ShapeFit and Full-Modeling approaches to extract cosmological information, and we fit the power spectrum monopole and quadrupole for $k\in[0.02,0.20]\,h\,\rm{Mpc}^{-1}$. The analysis is targeted on DESI- and Euclid-like survey with effective volume $V_{25}=25\,h^{-3}\rm{Gpc}^{3}$. Our key findings are summarized below:
Redshift uncertainty introduces diffusive effect in the clustering, leading to a scale-dependent suppression to the observed power spectrum. This smearing effect can be modeled by a Gaussian damping term [20]. We find that the EFT cournterterms, parameterized by $\alpha_{2}$ and $\alpha_{4}$, also capture this smearing effect, resulting in compressed and cosmological parameter estimates difference $5\%$ compared to the uncontaminated case. In contrast, redshift catastrophics cause an overall reduction on the power spectrum amplitude that approximately scales as $(1-f_{c})^{2}$. While this effect is negligible (difference $6\%$) for ELG-like catastrophics, it significantly biases the estimation of amplitude-related parameters for slitless-like errors, namely $df$ in ShapeFit and $\ln(10^{10}A_{s})$ in the Full-Modeling. In that case, we observe differences of $6\%{-}16\%$, corresponding to a $\sim 2.2\sigma$ shift relative to the clean case. To model the effect of redshift catastrophics, we use the correction factor $(1-f_{c})^{2}$ applied to the power spectrum in Eq.˜4.6. When freeing $f_{c}$, the extended model mitigates the bias in $\ln(10^{10}A_{s})$ to below $0.1\sigma$, but leads to $60\%$ degradation in its precision due to the strong degeneracy between $f_{c}$ and $\ln(10^{10}A_{s})$. Fixing $f_{c}$ to its expected value in the fit restores the constraining power, with only a modest $0.85\sigma$ deviation. This highlights the importance of accurately estimating the catastrophic failure rate $f_{c}$ and applying the appropriate correction model to ensure unbiased cosmological results in slitless spectroscopic surveys.
[…]
Future work should focus on refining the model of redshift errors, exploring higher-order statistics to complement the power spectrum analyses, using light-cone mocks and more observational data to validate and extend these findings. In particular, we plan to apply our framework to DESI DR2 analysis, which provides a large sample of repeat observations well-suited for characterizing redshift errors. Using this dataset, we aim to jointly model redshift uncertainty and redshift catastrophics for each tracer population, across a wide range of redshift, enabling a more realistic and tracer-specific treatment of redshift systematics for cosmological inference.
Acknowledgments
SH, AR, JY and JPK, acknowledge support from the Swiss National Science Foundation (SNF) ‘Cosmology with 3D Maps of the Universe” research grant 200020_207379. This research was essentially supported through computational resources provided by the High Performance Computing (HPC) facility of University of Geneva (UNIGE), especially for the Bamboo cluster.
Appendix A BAO fitting results
Catalogs clean dv-LRG dv-QSO-G dv-QSO-L dv-ELGcatas dv-slitless $\alpha_{\mathrm{iso}}$ $1.0031\pm 0.0045$ $1.0027\pm 0.0044$ $1.0026\pm 0.0045$ $1.0028\pm 0.0046$ $1.0034\pm 0.0046$ $1.0023\pm 0.0046$ $\alpha_{\mathrm{ap}}$ $1.001\pm 0.016$ $1.000\pm 0.015$ $0.996\pm 0.016$ $0.999\pm 0.015$ $1.000\pm 0.015$ $0.997\pm 0.016$ $\chi^{2}/\rm{dof}$ $4.51/33$ $5.45/33$ $6.25/33$ $5.89/33$ $4.67/33$ $5.23/33$
Table 4: Comparison of BAO fitting results for the clean and contaminated mock catalogs described in Table˜1. Table reports the mean posterior values with 68% confidence intervals. The $\alpha_{\mathrm{iso}}$ and $\alpha_{\mathrm{ap}}$ remain almost unbiased in the presence of redshift errors ($\Delta\alpha_{\mathrm{iso}}<0.18\,\sigma$ and $\Delta\alpha_{\mathrm{ap}}<0.32\,\sigma$). The corresponding best-fit $\chi^{2}$ over the degree of freedom (dof) is also listed for each catalog, showing an increase when redshift errors are included.
BAO serves as a standard ruler in cosmology and is generally considered robust against a wide range of systematic effects. In this work, we also test whether BAO measurements remain unbiased under the spectroscopic redshift errors. BAO analysis focus on constraining two dilation parameters, $\alpha_{\parallel}$ and $\alpha_{\perp}$, which are reparameterized into two scaling parameters, $\alpha_{\mathrm{iso}}$ and $\alpha_{\mathrm{ap}}$, respectively. These two parameters provide information about the isotropic BAO scale and AP effect. We adopt the BAO modeling framework implemented in DESI DR1 analyses [68] using the desilike package. The results are obtained by fitting the mean monopole and quadrupole of the two-point correlation function from 500 mock realizations without BAO reconstruction. The built covariance matrix are rescaled to an effective volume of $V_{25}=25\,h^{-3}\rm{Gpc}^{3}$.
Table˜4 summarizes the resulting BAO constraints. We find that the differences in $\alpha_{\mathrm{iso}}$ and $\alpha_{\mathrm{ap}}$ estimates remain small, with $\Delta\alpha_{\mathrm{iso}}<0.18\,\sigma$ and $\Delta\alpha_{\mathrm{ap}}<0.32\,\sigma$, and the corresponding error bars remain nearly unchanged. The best-fit $\chi^{2}$ values increase noticeably after including redshift errors, with up to a 40% increase observed in the dv-QSO-L sample. This indicates a decrease of the goodness of fit due to the redshift errors. But in general, the BAO measurements remain unbiased and robust against spectroscopic redshift errors at the current level of precision.
Appendix B $\chi^{2}$ results and contour plot for baseline analysis
Figure 9: Comparison of corresponding $\chi^{2}$ for Figure˜4 and Figure˜5. The best-fit $\chi^{2}$ values are normalized by the degrees of freedom, with results for ShapeFit (left panels) and Full Modeling (right panels). For each point, we fit the mean of 25 random mocks, with the covariance rescaled to $V_{25}$. The difference of $\chi^{2}$ is small between clean and contaminated mock catalogs, demonstrating that all fits achieve the same level of goodness.
Figure 10: Posterior distribution for nuissance parameters in our baseline EFT  fitting. Contours show the 68% and 95% confidence interval. Results are derived from fits to the mean power spectrum monopole and quadrupole of random selected 25 mocks using the covariance matrix rescaled to $V_{25}$. The negative shift of $\alpha_{2}$, $\alpha_{4}$ on their marginalized posterior are account for the effect of redshift uncertainty, especially for the QSO-like smearing cases.
Figure˜9 presents the distribution of best-fit $\chi^{2}$ values for the baseline analysis. The left two panels show results for the ShapeFit analysis, corresponding to the scatter plot in Figure˜4, while the right two panels show the Full-Modeling results for Figure˜5. From the $\chi^{2}$ distributions, we do not observe significant differences ($15\%$) or systematic shifts between the clean and contaminated mocks analyses. This indicates that the theoretical model provides a same level of good prediction with the observational data, even potential deviations in the parameter estimates may still arise under contamination.
Figure˜10 shows a representative posterior distribution for nuisance parameters in EFT of LSS framework, comparing results from the clean mocks and those contaminated by redshift uncertainty. We recall that the $b_{i}$ parameters correspond to the biased terms, $\alpha_{i}$ denote the EFT counterterms and $s_{n,i}$ model the stochastic contribution. The most noticeable shifts in the posteriors are observed for $\alpha_{2}$, $\alpha_{4}$, and $s_{n,2}$, especially for the QSO-like smearing cases. As explained in Section˜4.3.1, the EFT counterterms $\alpha_{2}$ and $\alpha_{4}$ are expected to absorb the damping effect in the power spectrum introduced by redshift uncertainty. This behavior is reflected as a negative shift in their estimated values, with the value increasing with the level of smearing. The parameter $s_{n,2}$ represents the second-order stochastic term, which scales as $(k\mu)^{2}$. Its variations likely reflect the stochastic fluctuations during the fitting, but we do not take them explicitly accounted for redshift uncertainty effect.